# POPGym Arcade: Parallel Pixelated POMDPs

### v5

@@ Abstract
We present the POPGym Arcade [^1] Code available at https://github.com/bolt-research/popgym_arcade, a collection of hardware-accelerated, pixel-based environments with shared observation and action spaces. Each environment includes fully and partially observable variants, enabling counterfactual studies on partial observability. We also introduce mathematical tools for analyzing policies under partial observability, which reveal how agents recall past information to make decisions. Our analysis shows (1) that controlling for partial observability is critical and (2) that agents with long-term memory learn brittle policies that struggle to generalize. Finally, we demonstrate that recurrent policies can be “poisoned” by old, out-of-distribution observations, with implications for sim-to-real transfer, imitation learning, and offline reinforcement learning.
@@ 1 Introduction
Real-world decision making often requires reasoning under partial observability, where important information is hidden from the agent. This is a challenge for reinforcement learning (RL), which excels in fully observable settings (Silver et al., 2016). Scalable, efficient benchmarks for partial observability remain scarce, and existing solutions often lack tools to analyze how agents adapt to incomplete information.
Benchmarks like the Arcade Learning Environment (ALE) (Bellemare et al., 2013) and MuJoCo (Todorov et al., 2012) have driven RL progress, but they provide full or near-full observability (Machado et al., 2018), limiting their utility for partially observable scenarios. Prior work also rarely investigates how agents internally compensate for missing information over time.
We introduce POPGym Arcade (Fig. 1), a pixel-based, hardware-accelerated benchmark inspired by POPGym (Morad et al., 2023a) and the ALE. Each environment includes “twins”: fully and partially observable variants with identical dynamics, enabling controlled studies of observability. We complement this with tools to visualize how agents store and recall information, addressing the opacity of decision making under partial observability. By combining hardware acceleration, environment twins, and pixel observations with our visualization tools, we perform efficient counterfactual studies on observability, in a high-dimensional and interpretable observation space.
@@
In this work, we contribute:
  • GPU-accelerated, Atari-style partially observable environments.
  • Fully observable twins for counterfactual studies on observability.
  • Tools to examine how agents store and recall information over time.
  • Evidence that solutions to partial observability introduce confounding factors.
  • Evidence that recurrent value functions do not generalize well.
MDP Variants
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
POMDP Variants
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Observations from our environment twins. All environments share a unifying observation, state, and action space, enabling counterfactual studies with observability as the sole independent variable.
2 Background
@@
In RL, we interact with an environment to maximize the discounted cumulative reward. A Markov Decision Process (MDP) (Sutton & Barto, 2018) is a tuple $\langle S,A,P,R,\gamma\rangle$, with the set of states $S$, actions $A$, and transition function $P:S\times A\mapsto\Delta(S)$, where $\Delta(S)$ is a distribution over $S$. MDPs also contain a reward function $R:S\times A\mapsto\mathbb{R}$ and discount factor $\gamma\in[0,1]$. At each timestep, we receive a state $s_{t}\in S$ and select an action $a_{t}\in A$ from a policy $a_{t}\sim\pi(\cdot|s_{t})$ that takes us to the next state following $P(s_{t+1}\mid s_{t},a_{t})$ while emitting a reward $r_{t}=R(s_{t},a_{t})$. We aim to learn the policy $\pi:S\mapsto\Delta(A)$ that maximizes the expected return $\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\mid s_{0}=s]$.
2.1 Partially Observable Markov Decision Processes
In many realistic scenarios, the agent does not have access to the full state of the environment, instead perceiving portions of the state through noisy or limited sensors. We call this problem setup a Partially Observable Markov Decision Process (POMDP). A POMDP is an extension of the MDP, consisting of tuple $\langle S,A,P,R,\gamma,\Omega,O\rangle$, with the set of observations $\Omega$ and the observation function $O:S\mapsto\Delta(\Omega)$. At each timestep, the agent can only access the current observation $o_{t}\sim O(s_{t})$. The observation is not guaranteed to be Markov, thus providing only partial observability of the state, and preventing naive application of standard RL approaches to POMDPs. Fortunately, the sequence of all prior observations and actions, known as the trajectory $\mathbf{x}_{t}=(o_{0},a_{0},o_{1},a_{1}\dots a_{t-1},o_{t})$, is itself Markov (Sutton & Barto, 2018). As such, an optimal policy should consider all prior observations and actions.
2.2 Memory and Recurrences
For POMDPs, the optimal policy must be a function of a variable length trajectory $\mathbf{x}_{t}$. For conciseness, we introduce the combined observation-action space $X=(\Omega,A)$ and represent the observation and previous action as $x_{t}=(o_{t},a_{t-1})$.
@@ 2.2 Memory and Recurrences
It is cumbersome to learn a policy over a variable size $\mathbf{x}_{t}$. We find it simpler to introduce a memory model $f$ that produces a fixed-size latent Markov state $\hat{s}\in\hat{S}$ from the trajectory $\mathbf{x}_{t}$. This confines the complexities associated with partial observability to $f$, allowing the policy to operate on a latent Markov state $\hat{s}$. More formally, a recurrent memory model $f:X\times H\mapsto\hat{S}\times H$ maintains a summary of the trajectory in the recurrent state $h\in H$ and outputs a latent Markov state $\hat{s}\in\hat{S}$ for the policy

### v6

# Investigating Memory in RL with POPGym Arcade

@@ Abstract
How should we analyze memory in deep RL? We introduce mathematical tools for fairly analyzing policies under partial observability and revealing how agents use memory to make decisions. To utilize these tools, we present POPGym Arcade [^1] Code available at https://github.com/bolt-research/popgym-arcade, a collection of Atari-inspired, hardware-accelerated, pixel-based environments sharing a single observation and action space. Each environment provides fully and partially observable variants, enabling counterfactual studies on observability. We find that controlled studies are necessary for fair comparisons, and identify a pathology where value functions smear credit over irrelevant history. With this pathology, we demonstrate how out-of-distribution scenarios can contaminate memory, perturbing the policy far into the future, with implications for sim-to-real transfer and offline RL.
@@ 1 Introduction
Real-world decision making often requires reasoning under partial observability, where important information is hidden from the agent. This is a challenge for reinforcement learning (RL), which excels in fully observable settings (Silver et al., 2016). To mitigate partial observability, we store and recall sensory information using some form of memory. We often evaluate deep memory models by comparing returns across partially observable tasks. However, deep RL experiments are sensitive to model configurations, observation size, task difficulty, and other confounding factors; especially when combined with memory (Mania et al., 2018; Pleines et al., 2025). This sensitivity makes it difficult to isolate the impact of memory on observability, fairly quantify memory performance, or understand failure modes (Jordan et al., 2024).
To address this gap in understanding, we introduce mathematical tools to isolate and study memory and observability, and to interpret how agents use memory. To utilize these tools, we introduce POPGym Arcade (Fig. 1), an Atari-style hardware-accelerated benchmark where every environment has fully observable and partially observable “twins” with identical spaces and dynamics. By combining twins with our memory analysis tools, we perform counterfactual studies on observability. Our findings motivate a secondary study investigating what memory models learn, where we uncover credit attribution failures.
MDP Environments
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
POMDP Environments
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Observations from our environment twins. All environments share a unifying observation and action space, enabling counterfactual studies with observability as the sole independent variable.
@@
  1. (C1)
    Tools to disentangle, measure, and interpret memory in an RL context.
  2. (C2)
    POPGym Arcade: a benchmark of paired MDP/POMDP twins for controlled studies.
  3. (C3)
    Evidence that memory introduces bias and confounds return-based comparisons.
  4. (C4)
    Discovery of a pathology where memory models incorrectly assign value to irrelevant history.
  5. (C5)
    Demonstration of recurrent state contamination, where C4 causes irrelevant past to corrupt a policy.
2 Preliminaries
Reinforcement Learning
@@ 2 Preliminaries
A Markov Decision Process (MDP) (Sutton & Barto, 2018) is a tuple $\mathcal{M}=\langle S,A,P,R,\gamma\rangle$, with the set of states $S$, actions $A$, and transition function $P:S\times A\mapsto\Delta(S)$, where $\Delta(S)$ is a distribution over $S$. MDPs also contain a reward function $R:S\times A\mapsto\Delta(\mathbb{R})$ and discount factor $\gamma\in[0,1)$. At each timestep, we receive a state $s_{t}\in S$ and select an action $a_{t}\in A$ from a policy $a_{t}\sim\pi(\cdot|s_{t})$ that takes us to the next state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$ emitting a reward $r_{t}\sim R(\cdot|s_{t},a_{t})$. The Markov property states that $s_{t}$ contains sufficient information for an optimal policy. In RL, we learn the policy $\pi:S\mapsto\Delta(A)$ that maximizes the expected return $J(\pi,\mathcal{M})=\mathbb{E}_{\pi,\mathcal{M}}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$ for $\mathcal{M}$.
POMDPs
In Partially Observable Markov Decision Processes (POMDPs), the agent cannot access environment state, instead perceiving the state through noisy or limited sensors. A POMDP $\mathcal{P}=\langle S,A,P,R,\gamma,\Omega,O\rangle$ extends the MDP with set of observations $\Omega$ and observation function $O:S\mapsto\Delta(\Omega)$. At each timestep, the agent indirectly accesses the state through a non-Markov observation $o_{t}\sim O(s_{t})$. Fortunately, the trajectory $\mathbf{x}_{t}=(o_{0},a_{0},o_{1},a_{1}\dots a_{t-1},o_{t})$, is itself Markov (Sutton & Barto, 2018). The Reward Memory Length (RML) determines how far back in time we must search to learn the optimal policy (Ni et al., 2024). An RML of $k$ expresses that the reward depends on only the latest $k$ elements of the trajectory $r_{t+1}\perp\!\!\!\!\perp\mathbf{x}_{t},a_{t}|\mathbf{x}_{t-k:t},a_{t}$. Our new RL objective depends on a memory model $f$ and POMDP $\mathcal{P}$: $J(f,\pi,\mathcal{P})=\mathbb{E}_{\pi,f,\mathcal{P}}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$ .
Memory and Recurrences
@@
A memory model $f$ produces a fixed-size latent Markov state $\hat{s}\in\hat{S}$ from the trajectory $\mathbf{x}_{t}$. This confines the complexities associated with partial observability to $f$, allowing the policy to operate on a latent Markov state $\hat{s}$. For conciseness, we introduce the joint observation-action space $X=(\Omega,A);x_{t}=(o_{t},a_{t-1})$. A recurrent memory model $f:X\times H\mapsto\hat{S}\times H$ maintains a summary of the trajectory in the recurrent state $h\in H$ and outputs a latent Markov state $\hat{s}\in\hat{S}$ for the policy